# How Does A Text Preprocessing Pipeline Affect Ontology Syntactic Matching?

### v2

@@ IX Conclusion
In this paper, we conduct a comprehensive study on the effect of the text preprocessing pipeline on syntactic OM. 8 OAEI track repositories with 49 distinct alignments are evaluated. Despite the importance of text preprocessing in syntactic OM, our experimental results indicate that the text preprocessing pipeline is currently ill-equipped to handle OM tasks. (1) We find that Phase 1 text preprocessing methods (Tokenisation and Normalisation) help with both matching completeness (i.e. Recall) and correctness (i.e. Precision). Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are less effective. They can improve matching completeness (i.e. Recall), but matching correctness (i.e. Precision) is relatively low. (2) We propose a novel context-based pipeline repair approach to repair the less effective Phase 2 text preprocessing methods. By using a reserved word set to reduce false positive samples detected, our approach outperforms the traditional text preprocessing pipeline, in particular, the matching correctness (i.e. Precision) and overall matching performance (i.e. F1 Score).
@@
The source code for this study is available from https://github.com/qzc438/ontology-nlp. The OAEI track repositories used in this study can be downloaded from OAEI MELT at https://dwslab.github.io/melt/track-repository (date accessed: 2024-06-01). The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.
@@ Acknowledgment
Appendix A Supplementary Material
Appendix B Extended Experiment Details
B-A Analysis of Text Preprocessing Pipeline
Fig. 121314, and 15 show the details of the experiment to compare the text preprocessing pipeline in syntactic OM.
Figure 12: Base without Preprocessing (B) vs. Tokenisation (T).
Figure 13: Tokenisation (T) vs. Normalisation (N).
Figure 14: Normalisation (N) vs. Stop Words Removal (R).
Figure 15: Stop Words Removal (R) vs. Stemming/Lemmatisation (S/L).
  1. (1)
    For Phase 1 text preprocessing methods (Tokenisation and Normalisation), most of the data points located above the equivalent line in Precision, Recall, and F1 Score indicate that they help syntactic OM.
  2. (2)
    For Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation): Some data points are located above the equivalent line in Recall, but the majority of data points located below the equivalent line in Precision and F1 Score indicate that they do not help syntactic OM.
Fig. 16 shows the details of the experiment to compare Stemming (S) and Lemmatisation (L). Most of the data points located above the L=S line in Precision and F1 Score indicate that using Lemmatisation is better than Stemming in syntactic OM (assuming that the post hoc correction is excluded).
Figure 16: Stemming (S) vs. Lemmatisation (L).
Fig. 1718, and 19 show the details of the experiment to compare different stemmers.
Figure 17: Porter Stemmer (SP) vs. Snowball Stemmer (SS).
Figure 18: Porter Stemmer (SP) vs. Lancaster Stemmer (SL).
Figure 19: Snowball Stemmer (SS) vs. Lancaster Stemmer (SL).
  1. (1)
    For Porter Stemmer vs. Snowball Stemmer, all data points located in the equivalent line indicate that there is no difference using Porter Stemmer and Snowball Stemmer in syntactic OM.
  2. (2)
    For Porter/Snowball Stemmer vs. Lancaster Stemmer, most of the data points the data points located above the equivalent line in Precision and F1 Score indicate that Porter/Snowball Stemmer is more effective than Lancaster Stemmer in syntactic OM.
Fig 20 shows the details of the experiment to compare Lemmatisation vs. Lemmatisation + POS Tagging. All data points located in the equivalent line indicate that using POS Tagging in Lemmatisation does not help syntactic OM.
Figure 20: Lemmatisation (L) vs. Lemmatisation + Part-Of-Speech Tagging (LT).
B-B Context-based Pipeline Repair
Fig. 2122232425, and 26 show the details of the experiment to consider the benefit of using context-based pipeline repair in Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation).
  1. (1)
    For Precision, most of the data points are located above the equivalent line, indicating that context-based pipeline repair significantly improves matching correctness.
  2. (2)
    For Recall, most of the data points are located in the equivalent line, and only a few data points are located below the equivalent line, indicating that context-based pipeline repair slightly reduces matching completeness.
  3. (3)
    For F1 Score, most of the data points are located above the equivalent line, indicating that context-based pipeline repair also improves overall matching performance.
  4. (4)
    Experimentally, the wider ellipse around the data points indicates that the matching performance improvement ranking in Phase 2 text preprocessing methods is Stemming (S) > Lemmatisation (L) > Stop Words Removal (R).
  5. (5)
    Experimentally, the wider ellipse around the data points indicates that the matching performance improvement ranking in different stemmers is Lancaster Stemmer (SL) > Porter Stemmer (SP) = Snowball Stemmer (SS).
  6. (6)
    Experimentally, the same ellipse around the data points indicates that the matching performance improvement ranking in lemmatisation is Lemmatisation (L) = Lemmatisation + POS Tagging (LT).
Figure 21: Comparison of using and without using context-based repair in Stop Words Removal (R).
Figure 22: Comparison of using and without using context-based repair in Porter Stemmer (SP).
Figure 23: Comparison of using and without using context-based repair in Snowball Stemmer (SS).
Figure 24: Comparison of using and without using context-based repair in Lancaster Stemmer (SL).
Figure 25: Comparison of using and without using context-based repair in Lemmatisation (L).
Figure 26: Comparison of using and without using context-based repair in Lemmatisation + POS Tagging (LT).

### v3

@@ IX Conclusion
In this paper, we conduct a comprehensive study on the effect of the text preprocessing pipeline on syntactic OM. 8 OAEI track repositories with 49 distinct alignments are evaluated. Despite the importance of text preprocessing in syntactic OM, our experimental results indicate that the text preprocessing pipeline is currently ill-equipped to handle OM tasks. (1) We find that Phase 1 text preprocessing methods (Tokenisation and Normalisation) help with both matching completeness (i.e. Recall) and correctness (i.e. Precision). Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are less effective. They can improve matching completeness (i.e. Recall), but matching correctness (i.e. Precision) is relatively low. (2) We propose a novel context-based pipeline repair approach to repair the less effective Phase 2 text preprocessing methods. By using a reserved word set to reduce false positive samples detected, our approach outperforms the traditional text preprocessing pipeline, in particular, the matching correctness (i.e. Precision) and overall matching performance (i.e. F1 Score). (3) We also argue the traditional text preprocessing pipeline is not obsolete in the era of LLMs, but an integral part for LLMs used in the OM tasks.
@@
The source code for this study is available from https://github.com/qzc438-research/ontology-nlp. The OAEI track repositories used in this study can be downloaded from OAEI MELT at https://dwslab.github.io/melt/track-repository (date accessed: 2024-06-01). The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.